Suppose that Noah and Naomi have a monopoly in the garden bench market. Their weekly demand is D(P) = 500 - 4P. What is the inverse demand function PLEASE SHOW YOUR WORK
Answer #1

Answer

D(p)=500-4P

the inverse demand function is a function of quantity with price

D(p)=qauntity=Q

Q=500-4P

Q+4P=500

4P=500-Q

P=125-0.25Q

the inverse demand curve is P=125-0.25Q